OpenStudy (anonymous):

A ball is dropped from a height of 16 feet. Each time it drops h feet, it rebounds 0.81h feet. Find the total distance traveled by the ball.

OpenStudy (anonymous):

I got 152.421 feet, but I just want to make sure I did it right

OpenStudy (anonymous):

16 + 2(16*.81 + 16*.81*.81+...) It's a GP

OpenStudy (amistre64):

an = .81 a(n-1) an = a0 * .81^n or some such

OpenStudy (amistre64):

geometric progression, yes

OpenStudy (anonymous):

Lol, thanks I got confused over its name. Geometric for an algebraic series sounds weird. Does anyone here knows why it's called Geometric Progression?

OpenStudy (anonymous):

\[16 + 32(\frac{81}{100})\sum_{0}^{\infty}(\frac{81}{100})\] right?

OpenStudy (amistre64):

geometric stuff tends to be multiplications; like finding the area of a square

OpenStudy (anonymous):

\[16 + 32(\frac{81}{100})\sum_{0}^{\infty}(\frac{81}{100})^n\] I mean

OpenStudy (amistre64):

16 + 16*.81 + 16*.81 + 16*.81^2+ ... 16 + 2(16*.81 + 16*.81^2+ 16*.81^3+...) \[16+2(16\frac{1-.81^{inf}}{1-.81})\]

OpenStudy (amistre64):

the top tends to bow out to 1 i beleive

OpenStudy (amistre64):

\[16+\frac{32}{.19}\]maybe?

OpenStudy (turingtest):

I was thinking\[16+32\sum_{n=1}^{\infty}(0.81)^n=16+32[-1+\sum_{n=0}^{\infty}(0.81)^n]\]\[=-16+\frac{32}{1-0.81}\]because we have to strip out a term, no?

OpenStudy (amistre64):

might have to factor out the .81 as well to get the first term to 1

OpenStudy (anonymous):

amistre your answer is 32 off from mine, and turingtest, that's what i got :)

OpenStudy (amistre64):

\[16+\frac{32*.81}{.19}\]is my gut feeling in the end :)

OpenStudy (amistre64):

152.42.... yeah :)

OpenStudy (amistre64):

\[(1+r+r^2+r^3+...)=\frac{1}{1-r}\] since we had a 16*.81 at the start we need to factor it out to get this format; and then add in the original 16 that it fell from